OpenStudy (anonymous):

Le f be a function such that f(x+y)=f(x)+f(y) for all real x,y. Suppose lim as x goes to 0=L . Show that there exists a such that f(x)=ax for all x

OpenStudy (zarkon):

Is L supposed to be any number??

OpenStudy (zarkon):

Wouldn't you want the limit as x goes to 0 to be 0?

OpenStudy (anonymous):

yes sorry L=0

OpenStudy (anonymous):

it was the first part of the question to show, I also showed that \[lim_{x \rightarrow c}f(x)=f(c)\]

OpenStudy (zarkon):

what is your guess for the value of a?

OpenStudy (anonymous):

1 would be my guess

OpenStudy (zarkon):

you can you simplify f(2) as \[f(2)=f(1+1)=f(1)+f(1)=f(1)\cdot 2\] what would f(3),f(4),... and so on be?

OpenStudy (zarkon):

*you can simplify...

OpenStudy (anonymous):

oh ok let me try

OpenStudy (anonymous):

ok so I can get \[f(x)=f(x/2)*2 \] but then how do I jump to f(x)=ax?

OpenStudy (zarkon):

and how did you get that?

OpenStudy (zarkon):

can you answer my question from above..what is f(3),f(4),...?

OpenStudy (anonymous):

ah yes ok I get it. \[ f(3)=f(2)+f(1)=3*f(1) , f(4)=4f(1)....\] therefore a=f(1) ? It works nicely for the naturals but for all real?

OpenStudy (zarkon):

...we will build up to the reals..

OpenStudy (zarkon):

can you show that f(-n)=-f(n)

OpenStudy (zarkon):

where n is a natural number

OpenStudy (anonymous):

should I start by rewriting \[f(-n) =f(n-2n)\] or would does it follow from the addition property ?

OpenStudy (zarkon):

hint: 0=1+(-1)

OpenStudy (anonymous):

ok so \[f(-n)=f(-n)+f(n)-f(n) =f(-n+n)-f(n)=f(0)-f(n)=-f(n)\]

OpenStudy (zarkon):

ok that works...I was thinking \[0=f(0)=f(n+(-n))=f(n)+f(-n) \text{ so }f(-n)=-f(n)\]

OpenStudy (zarkon):

now you can do negative integers \[f(-n)=-f(n)=-f(1)n\]

OpenStudy (zarkon):

now can you show \[f\left(\frac{1}{m}\right)=f(1)\frac{1}{m}\]

OpenStudy (zarkon):

where \(m\) is a natural number

OpenStudy (anonymous):

should I write f(1/m) as a sum of other terms?

OpenStudy (zarkon):

\[\begin{array}{lc} 1= & \underbrace{\frac{1}{m}+\frac{1}{m}+\cdots +\frac{1}{m}}\\ & m \text{ terms} \\ \end{array}\]

OpenStudy (anonymous):

ok so\[f(1)=mf(1/m) \rightarrow\frac{1}{m}f(1)=f(1/m)\]

OpenStudy (zarkon):

why does \(f(1)=mf(1/m)\)? you need to show that

OpenStudy (zarkon):

\[f(1)=f\left(\frac{1}{m}+\frac{1}{m}+\cdots+\frac{1}{m}\right)=f\left(\frac{1}{m}\right)m\]

OpenStudy (zarkon):

I assume you used my hint

OpenStudy (anonymous):

\[f(1)=f(1/m+1/m...1/m) =f(1/m)+f(1/m)+...+f(1/m) \text{ m times} \] \[ f(1)=mf(1/m) \text{ dividing by m } f(1)*\frac{1}{m}=f(1/m) \]

OpenStudy (zarkon):

ok.. almost done.... show \[f\left(\frac{n}{m}\right)=f(1)\frac{n}{m}\] where \(n,m\in\mathbb{N},m\neq 0\)

OpenStudy (zarkon):

we could let \(n\in\mathbb{Z}\)

OpenStudy (zarkon):

or \(m\in\mathbb{Z}\)

OpenStudy (anonymous):

os so \[ f(\frac{n}{m})=f(1/m)+f(1/m)+...f(1/m) \text{ n times} \] \[f(n/m)=f(1/m)n \text{ using the previous result we have} f(1/m)=f(1)\frac{1}{m} \] \[f(n/m)=f(1)\frac{n}{m}\]

OpenStudy (zarkon):

\(m\) obviously not zero

OpenStudy (zarkon):

ok

OpenStudy (zarkon):

now you proved that \[\lim_{x\to c}f(x)=f(c)\] correct?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

ok...let \(x\in\mathbb{R}\) let \[\{x_n\}\subset\mathbb{Q}\] where \[x_n\to x\] (we can do this since the rational numbers are dense in the reals)

OpenStudy (zarkon):

then \[\lim_{n\to\infty}f(x_n)=f(x)\]

OpenStudy (zarkon):

but \[f(x_n)=f(1)x_n\] by our previous results ...

OpenStudy (anonymous):

ok so you are taking Xn as a subest of the rationals , as subsequence that goes to x? and then since \[ f(x_n)=f(1)x_n \] and because of the limit of f(Xn)=f(x) it follows that f(x)=f(1)x? hence a=f(1)

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

wow thank you so much! thanks for being so patient !

OpenStudy (zarkon):

no problem